A rectangle has a perimeter of 60 inches and lenth of 22 inches. What is the width of the rectangle?
Answer by checkley77(12844)

L=22
2L+2W=60
2*22+2W=60
44+2W=60
2W=60-44
2W=16
W=16/2
W=8 ANS. FOR THE WIDTH.
PROOF:
2*22+2*8=60
44+16=60
60=60